OpenStudy (anonymous):

How do I solve tan(sin^-1(x/2))?

myininaya (myininaya):

\[\text{ Let } u=\sin^{-1}(\frac{x}{2}) => \sin(u)=\frac{x}{2} (=\frac{opp}{hyp})\] Label a right triangle with the opp side having measurement x and the hyp having measurement 2 Then use that triangle to find tan(u) in terms of x

OpenStudy (anonymous):

Thank you!!